OpenStudy (anonymous):

Solve the equation by completing the square: X^2+5x-9=0

OpenStudy (uri):

@Mimo1997 Quadratic equation? Do you know the formula?

OpenStudy (anonymous):

Yea but i think I'm getting it wrong.

OpenStudy (anonymous):

It's a(x-d)^2+e=0 where d= b/2a and e=b^2/4a @uri